# Understanding LLM Agent Behaviours via Game Theory: Strategy Recognition, Biases and Multi-Agent Dynamics

### v1

@@ 4.1 Summary
This study introduces an integrated framework for understanding LLM agent behaviours via game-theoretic benchmarks, strategy recognition, and bias analysis in both dyadic and multi-agent settings. Building on FAIRGAME, we design controlled repeated-game environments that probe how LLMs respond to incentives and social dilemmas: a payoff-scaled Prisoner’s Dilemma to manipulate the stakes of cooperation without altering the underlying game, and a three-player Public Goods Game with configurable personalities, languages, and payoff parameters to elicit rich multi-agent dynamics such as free-riding, coordination, and coalition-like behaviour. On top of these simulations, we develop a machine learning-based intent recognition pipeline that encodes game trajectories into state-action sequences, trains classifiers on canonical strategies (ALLC, ALLD, TFT, WSLS) under execution noise, and then applies these models to FAIRGAME logs to infer the latent strategies and systematic biases of LLM agents across models, languages, and roles.
@@
Moreover, we conduct a systematic study of the interaction between language and cognition by evaluating the performance of agents across diverse language prompts and with different payoff matrix coefficients. This allows us to observe how language formulations affect economic rationality across varying levels of payoff matrix importance, revealing profound asymmetries between languages, where agents demonstrating optimal strategies in high-resource languages may exhibit suboptimal or anomalous behavior such as a failure to synchronize in avoiding penalties when prompted in low-resource languages.
@@
By analyzing LLM behaviors and predicting their induced strategies using a machine-learning classifier in a repeated Prisoner’s Dilemma setting, we observe that different large language models systematically prefer different strategic profiles. Claude tends to be the most cooperative, while Llama3-405B-Instruct frequently defaults to an always-defect pattern. Mistral-Large shows stronger preferences for TFT and WSLS, whereas GPT-4o predominantly favors WSLS.
In addition to model-specific tendencies, we also find clear cross-linguistic effects: the preferred strategies shift depending on the language of the prompt. This may reflect cultural or dataset-driven biases embedded in the training process. For instance, Arabic and Vietnamese prompts often elicit more defect-oriented behaviors, such as a higher likelihood of adopting the ALLD strategy.
4.2 Limitations and future works
4.2.1 Limitations
Our study presents several methodological constraints that limit the generalizability of findings. First, the ten-round game horizon across both Prisoner’s Dilemma and Public Goods Game experiments, while sufficient to observe initial strategic patterns and end-game dynamics, may be too short to capture sophisticated long-term behaviours such as reputation-building, conditional reciprocity, or forgiveness strategies that require extended interaction histories to stabilize. Prior work on human subjects shows that experiments under 25 rounds primarily capture the learning phase, not stabilized strategic behavior, meaning that short sequences risk misrepresenting the true strategy dynamics [inferenceStrategies] Second, our linguistic coverage is restricted to English and Vietnamese, preventing broader conclusions about how linguistic-cultural priming operates across diverse language families with different grammatical structures, collectivist-individualist orientations, or varying representation densities in LLM training corpora. Third, our experimental design examines only two game-theoretic settings with limited parametric variation-the Prisoner’s Dilemma explores payoff magnitude scaling but maintains symmetric two-player structure, while the Public Goods Game uses a fixed group size of three agents, which represents the minimal configuration for non-dyadic effects but precludes examination of larger-group phenomena such as diffusion of responsibility or coalition formation. Fourth, the machine learning-based intent recognition pipeline focuses exclusively on four canonical strategies (ALLC, ALLD, TFT, WSLS) derived from classical game theory, and our high-confidence filtering approach prioritizes interpretability while necessarily excluding behavioural patterns that may represent emergent hybrid strategies not captured by this taxonomy. Finally, the absence of parallel human behavioural experiments under matched conditions prevents rigorous assessment of whether the observed cross-lingual cooperation gaps, model-specific prosocial biases, and strategic patterns genuinely reflect human-like reasoning or constitute artifacts of model-specific alignment procedures, limiting our ability to evaluate the ecological validity of LLM agents in real-world multi-agent scenarios.
4.2.2 Future work
Future work will focus on advancing strategy identification to understand how LLM agents behave in repeated social dilemmas and whether these behaviours differ from those of humans, using game-playing trajectories as the basis for comparison. We will collect substantially more LLM data by extending the original FAIRGAME 10-round interactions into long 100+-round games, increasing observations for downstream analysis and expanding contextual information by incorporating both action histories and Chain-of-Thought (CoT) [NEURIPS2022_9d560961] traces into strategy inference. With this larger dataset, we will explore more efficient strategy-identification techniques-such as clustering on extended trajectories, clustering on CoT sequences combined with actions, supervised models beyond logistic regression and LSTMs, and unsupervised methods including Hidden Markov Models [inferenceStrategies]-to support both classification and discovery of emergent strategies. We will benchmark these inferred strategies against human behavioural datasets to analyse similarities and divergences between LLM and human strategic behaviour, and pursue new directions such as multi-agent (n-player) games beyond two-player IPD/PGG. Finally, we will extend the FAIRGAME setup by experimenting with varied prompt templates, adding more languages beyond English and Vietnamese, enabling dynamic-importance or continual-learning configurations, mixing different games within multi-round sessions to test whether strategy shifts occur, and optionally allowing inter-agent communication to examine how environmental and linguistic factors shape LLM strategy formation.
@@ A.2 Public Goods Game implementation
This section provides full pseudocode for the multi-agent Public Goods Game (PGG) implementation used in our FAIRGAME extension. The algorithms mirror the original FAIRGAME control flow (game creation, execution loop, and per-round interaction), but are adapted to (i) instantiate three LLM-based agents per game, (ii) attach a dynamic public-goods payoff module, and (iii) handle vector-valued histories of joint actions and payoffs.
@@
 // dynamic public-goods payoff
@@ A.3 Intent Recognition via Machine Learning
As illustrated in Figure A1, English interactions were characterized by a hyper-competitive baseline, exhibiting the highest density of Always Defect (AllD) strategies and the lowest rates of adaptive cooperation. This behavior likely reflects the dominance of game-theoretic and individualistic maximizing narratives in the Anglo-centric training corpus. Conversely, Vietnamese prompts elicited the highest frequency of unconditional cooperation (AllC), consistent with the hypothesis that the model retrieves collectivist or community-oriented norms associated with the language.
@@
Figure A1: The Language Effect. (Left) Average payoffs achieved by agents across linguistic settings. (Right) Strategy distribution revealing cultural heterogeneity: English prompts drive competitive defection (AllD), Chinese prompts favor reciprocal strategies (TFT), while French prompts encourage adaptive, reinforcement-learning-style behaviors (WSLS), distinct from the high unconditional cooperation (AllC) observed in Vietnamese.
@@
Finally, we analyzed the impact of agent ordering on strategic adoption, revealing a distinct "Positional Bias" as illustrated in Figure A2. The empirical data indicates a significant divergence in behavior contingent upon role assignment: Agents designated as "Agent 1" (positioned initially in the system prompt) exhibited a marked propensity for aggressive, non-cooperative strategies, predominantly converging on Always Defect (ALLD). Conversely, "Agent 2" displayed a broader, more reactive strategic repertoire, often defaulting to cooperative behaviors (AllC) or conditional strategies.
We hypothesize that this asymmetry is an artifact of the autoregressive nature of Large Language Models combined with the "Primacy Effect." The sequential primacy of the first-mentioned entity in the prompt appears to be encoded by the model as an implicit cue for higher status or a "first-mover advantage." This "Implicit Hierarchy" underscores a critical methodological consideration: the ordering of agents in multi-agent simulations is not a neutral variable and can systematically skew negotiation dynamics and collective outcomes.
Figure A2: Role Asymmetry in Strategy Selection. A comparative analysis of strategy distribution between Agent 1 and Agent 2. Agent 1 (left) demonstrates a dominant preference for defecting strategies (AllD), whereas Agent 2 (right) exhibits a higher frequency of cooperative behaviors (AllC), suggesting an implicit hierarchy derived from prompt ordering.

### v2

@@ 6 Conclusion
We have introduced an integrated framework for understanding LLM agent behaviours via game-theoretic benchmarks, strategy recognition, and bias analysis in both dyadic and multi-agent settings. Building on FAIRGAME, we design controlled repeated-game environments to examine how LLMs respond to varying incentives and social dilemmas: a payoff-scaled Prisoner’s Dilemma to manipulate the stakes of cooperation without altering the underlying game, and a three-player Public Goods Game with configurable personalities, languages, and payoff parameters to elicit rich multi-agent dynamics such as free-riding, coordination, and coalition-like behaviour. On top of these simulations, we develop a machine learning-based intent recognition pipeline that encodes game trajectories into state-action sequences, trains classifiers on canonical strategies (ALLC, ALLD, TFT, WSLS) under execution noise, and then applies these models to FAIRGAME logs to infer the latent strategies and systematic biases of LLM agents across models, languages, and roles.
@@
Moreover, we conduct a systematic study of the interaction between language and cognition by evaluating the performance of agents across diverse language prompts and with different payoff matrix coefficients. This allows us to observe how language formulations affect economic rationality across varying levels of payoff matrix importance, revealing profound asymmetries between languages, where agents demonstrating optimal strategies in high-resource languages may exhibit suboptimal or anomalous behaviour such as a failure to synchronize in avoiding penalties when prompted in low-resource languages.
@@
By analyzing LLM behaviours and predicting their induced strategies using a machine-learning classifier in a repeated Prisoner’s Dilemma setting, we observe that different LLM systematically prefer different strategic profiles. Claude tends to be the most cooperative, while Llama3-405B-Instruct frequently defaults to an always-defect pattern. Mistral-Large shows stronger preferences for TFT and WSLS, whereas GPT-4o predominantly favors WSLS.
In addition to model-specific tendencies, we also find clear cross-linguistic effects: the preferred strategies shift depending on the language of the prompt. This may reflect cultural or dataset-driven biases embedded in the training process. For instance, Arabic and Vietnamese prompts often elicit more defect-oriented behaviours, such as a higher likelihood of adopting the ALLD strategy.
@@ A.2 Public Goods Game implementation
This section provides full pseudocode for the multi-agent Public Goods Game (PGG) implementation used in our FAIRGAME extension. The algorithms mirror the original FAIRGAME control flow (game creation, execution loop, and per-round interaction), but are adapted to (i) instantiate three LLM-based agents per game, (ii) attach a dynamic public goods payoff module, and (iii) handle vector-valued histories of joint actions and payoffs.
@@
 // dynamic public goods payoff
@@ A.3 Intent Recognition via Machine Learning
As illustrated in Figure A1, English interactions were characterized by a hyper-competitive baseline, exhibiting the highest density of Always Defect (AllD) strategies and the lowest rates of adaptive cooperation. This behaviour likely reflects the dominance of game-theoretic and individualistic maximizing narratives in the Anglo-centric training corpus. Conversely, Vietnamese prompts elicited the highest frequency of unconditional cooperation (AllC), consistent with the hypothesis that the model retrieves collectivist or community-oriented norms associated with the language.
@@
Figure A1: The Language Effect. (Left) Average payoffs achieved by agents across linguistic settings. (Right) Strategy distribution revealing cultural heterogeneity: English prompts drive competitive defection (AllD), Chinese prompts favor reciprocal strategies (TFT), while French prompts encourage adaptive, reinforcement-learning-style behaviours (WSLS), distinct from the high unconditional cooperation (AllC) observed in Vietnamese.
@@
Finally, we analysed the impact of agent ordering on strategic adoption, revealing a distinct "Positional Bias" as illustrated in Figure A2. The empirical data indicates a significant divergence in behaviour contingent upon role assignment: Agents designated as "Agent 1" (positioned initially in the system prompt) exhibited a marked propensity for aggressive, non-cooperative strategies, predominantly converging on Always Defect (ALLD). Conversely, "Agent 2" displayed a broader, more reactive strategic repertoire, often defaulting to cooperative behaviours (AllC) or conditional strategies.
We hypothesize that this asymmetry is an artifact of the autoregressive nature of LLMs combined with the "Primacy Effect." The sequential primacy of the first-mentioned entity in the prompt appears to be encoded by the model as an implicit cue for higher status or a "first-mover advantage." This "Implicit Hierarchy" underscores a critical methodological consideration: the ordering of agents in multi-agent simulations is not a neutral variable and can systematically skew negotiation dynamics and collective outcomes.
Figure A2: Role Asymmetry in Strategy Selection. A comparative analysis of strategy distribution between Agent 1 and Agent 2. Agent 1 (left) demonstrates a dominant preference for defecting strategies (AllD), whereas Agent 2 (right) exhibits a higher frequency of cooperative behaviours (AllC), suggesting an implicit hierarchy derived from prompt ordering.